OpenStudy (anonymous):

Stephanie is 3 yrs older than andrew. Twelve years from now the sum of their ages will be 30 yrs more than Stephanie's age at present. what expression represents the sum of their ages in 12 years?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Let x = Stephanie's age. Then x - 3 = Andrew's age.

OpenStudy (anonymous):

30[(x+3)+7] = age of mary in 7 years?

OpenStudy (skullpatrol):

Who is "mary"?

OpenStudy (anonymous):

i mean lol stephanie

OpenStudy (skullpatrol):

Why multiply by 30?

OpenStudy (anonymous):

sorry I'm very confused right now

OpenStudy (skullpatrol):

No problem, I will help you think through it...

OpenStudy (skullpatrol):

Let x = Stephanie's age. Then x - 3 = Andrew's age. Twelve years from now the sum of their ages will be 30 yrs more than Stephanie's age at present.

OpenStudy (skullpatrol):

How old will they be in 12 years?

OpenStudy (anonymous):

x+3= 12 (stephanie)?

OpenStudy (skullpatrol):

In 12 years: x + 12= Stephanie's age. Then x - 3 + 12 = Andrew's age.

OpenStudy (skullpatrol):

Please don't give the answer away.

OpenStudy (anonymous):

still confused here ;'(

OpenStudy (anonymous):

sorry

OpenStudy (skullpatrol):

Do you understand Let x = Stephanie's age. Then x - 3 = Andrew's age.

OpenStudy (anonymous):

yes i understand that part

OpenStudy (tammy18):

Andrew= x years old Stephenie = x + 3 Sum of their age now = x+x+ 3 = 2x + 3 Sum of their ages in twelve years 2x +3 + 12 = 30+x+3 x=18

OpenStudy (skullpatrol):

How about? In 12 years: x + 12= Stephanie's age. Then x - 3 + 12 = Andrew's age.

OpenStudy (anonymous):

Ahh I see

OpenStudy (skullpatrol):

Good :)

OpenStudy (skullpatrol):

Twelve years from now the sum of their ages will be 30 yrs more than Stephanie's age at present.

OpenStudy (skullpatrol):

@bless What is the equation that represents the relationships among the numbers in the problem?

OpenStudy (anonymous):

x+3+12 x+12

OpenStudy (skullpatrol):

Twelve years from now the sum of their ages: x + 12= Stephanie's age. x - 3 + 12 = Andrew's age.

OpenStudy (anonymous):

i got it :) thank u

OpenStudy (skullpatrol):

x + 12 + x - 3 + 12 = 30 + x

OpenStudy (skullpatrol):

Thanks for trying to learn.

OpenStudy (anonymous):

:D